The RMS value of e is 5V and RMS value of voltage drop across L is 4V. What is the RMS value of voltage across R?

Grade:12

1 Answers

Vikas TU
7 years ago
In R-L ac circuit,
Vnet = root(Vl^2 + Vr^2)
      5 = root(4^2 + Vr^2)
Vr = root(25 – 16) = > root(9) = > 3 volts.